$SPAD/src/input intmix2.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool intmix2.output
)set message test on
)set message auto off
)clear all
--S 1 of 4
(x + 1) / (x * (x + log x)**(3/2))
--R
--R
--R                    x + 1
--R   (1)  ----------------------------
--R                     2  +----------+
--R        (x log(x) + x )\|log(x) + x
--R                                                    Type: Expression(Integer)
--E 1
--S 2 of 4
integrate(%, x)
--R
--R
--R            +----------+
--R          2\|log(x) + x
--R   (2)  - --------------
--R            log(x) + x
--R                                         Type: Union(Expression(Integer),...)
--E 2

———-

This one does not have an elementary integral — * —

--S 3 of 4
log(1 + exp x)**(1/3) / (1 + log(1 + exp x))
--R
--R
--R          +------------+
--R         3|      x
--R         \|log(%e  + 1)
--R   (3)  ----------------

--R              x
--R        log(%e  + 1) + 1
--R                                                    Type: Expression(Integer)
--E 3
--S 4 of 4
integrate(%, x)
--R
--R
--R               +-------------+
--R           x  3|      %P
--R         ++   \|log(%e   + 1)
--R   (4)   |   ----------------- d%P
--R        ++         %P
--R             log(%e   + 1) + 1
--R                                         Type: Union(Expression(Integer),...)
--E 4
)spool
)lisp (bye)

———-

References^{[1] nothing}

5